OpenStudy (anonymous):

Find d2y/dx2 in terms of x and y. y^4 = x^5 d2y/dx2 =

OpenStudy (shamil98):

Find dy/dx first. Then differentiate it again to get d2y/dx2

OpenStudy (anonymous):

4y^3y'=5x^4 y'=5x^4/4y^3=5x/4y*xy/xy

OpenStudy (anonymous):

5y/4x*x^5/y^4= 5y/4x

OpenStudy (anonymous):

y"= 4x(5y')-5x(4)/16x^2

OpenStudy (anonymous):

i dont know how can i finish it :/

OpenStudy (shamil98):

y^4 = x^4 dy/dx 4y^3y' = 5x^4 y' = (5x^4)/(4y^3) \[y'' = \frac{ 20x^34y^3 - 5x^412y^2y' }{ (4y^3)^2 }\] Plug in y' into this and simplify it to get y'' :) Simplify if needed.

OpenStudy (shamil98):

\[\large y'' = \frac{ 20x^34y^3 - (5x^412y^2*\frac{ 5x^4 }{ 4y^3 }) }{ 16y^6 }\] Just simplify it out now.

OpenStudy (anonymous):

40y/16x^2

OpenStudy (anonymous):

@shamil98 right

OpenStudy (anonymous):

i dont know how can i do the top @shamil98